OpenStudy (anonymous):

Find the value of 10 times the square root of 2 over the square root of 5..

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

Hello

OpenStudy (misty1212):

\[\frac{10\sqrt{2}}{\sqrt{5}}\]?

OpenStudy (anonymous):

Yes

OpenStudy (misty1212):

they probably mean "rationalize the denominator" multiply by \[\frac{\sqrt5}{\sqrt5}\]

OpenStudy (misty1212):

\[\frac{10\sqrt{2}}{\sqrt{5}}\times\frac{\sqrt5}{\sqrt5}=\frac{10\sqrt{10}}{5}\]\

OpenStudy (misty1212):

then cancel

OpenStudy (anonymous):

5?